# ULog: Unsupervised Log Parsing with Large Language Models through Log Contrastive Units

### v1

@@ 3.6. Parallelization
In previous sections, we have discussed the building components of ULog. In practice, scalability is a crucial issue when parsing vast amounts of log messages (Wang et al., 2022). Therefore, we introduce the parallelization mechanism of ULog to improve the parsing speed.
As mentioned in §3.2, hierarchical sharding produces mutually exclusive log buckets at two levels. Consequently, logs in different buckets can be processed in a completely parallel fashion. However, since bucket updating is performed among all the buckets of the same length at the first level. Simply allocating buckets at the second level to produce template and update buckets can lead to template conflict, i.e., different templates being assigned to the same log. To address the problem, the allocation is conducted at the first level. Specifically, we aim to allocate $n$ executors for all first-level log buckets. Within each first-level log bucket, ULog iteratively selects the largest second-level bucket, selects an LCU, queries the LLM to obtain the template, and updates buckets under this level.
@@ 4. Experiment Setup
We conduct extensive experiments to evaluate ULog by answering the following research questions (RQs):
  • RQ1: How effective is ULog?
  • RQ2: How do different settings affect ULog?
  • RQ3: How efficient is ULog?
@@ 4.1. Datasets
We evaluate ULog on Loghub-2.0 (He et al., 2020; Jiang et al., 2023b), a collection of large-scale benchmark datasets for log parsing provided by LogPAI team (Zhu et al., 2019). Loghub-2.0 is annotated with ground-truth templates for 14 diverse log datasets, which covers a variety of systems such as distributed systems, supercomputer systems, and server-side applications. Compared with the original Loghub (He et al., 2020), Loghub-2.0 is equipped with 3×3$\times$3 × templates and 1,800×1,800$\times$1 , 800 × log messages on average, which can support more comprehensive evaluation on accuracy and efficiency. On average, each dataset in Loghub-2.0 comprises 3.6 million log messages. In total, the collection contains approximately 3,500 unique log templates.
@@ 4.2. Baselines
We compare ULog with seven open-sourced state-of-the-art log parsers, including four unsupervised syntax-based parsers and three label-required semantic-based parsers. For unsupervised label-free parsers, we adopt three syntax-based methods and one LLM-based method. The syntax-based methods include AEL (Jiang et al., 2008), Drain (He et al., 2017), and Brain (Yu et al., 2023), chosen for their superior performance compared to other syntax-based methods (Zhu et al., 2019; Khan et al., 2022; Jiang et al., 2023b). As there are currently no LLM-based log parsers proposed for label-free parsing, we adopt a label-free variant of LILAC (Jiang et al., 2023a) (LILAC w/o ICL) by removing the in-context learning (ICL) module. In terms of label-required log parsers, we select three state-of-the-art semantic-based log parsers: UniParser (Liu et al., 2022), LogPPT (Le and Zhang, 2023), and LILAC (Jiang et al., 2023a). UniParser trains a long short-term memory (LSTM) model on labelled log data for log parsing. LogPPT uses labelled log data to perform prompt-based fine-tuning based on RoBERTa (Liu et al., 2019a). LILAC samples similar logs from a small set of labelled logs (e.g., 32 logs) for each log, utilizing an ICL paradigm for LLMs to parse logs.
@@ 4.3. Evaluation Metrics
Following previous works (Zhu et al., 2019; Khan et al., 2022; Jiang et al., 2023b), we evaluate ULog with the following four metrics.
@@ 4.4. Environment and Implementation
We conduct all the experiments on a Ubuntu 20.04.4 LTS server with 256RAM and an NVIDIA A100 40G GPU. The LLM used in ULog is ChatGPT (gpt-3.5-turbo-0613) due to its popularity in recent log analysis studies (Xu et al., 2024; Jiang et al., 2023a; Ma et al., 2024). We invoke the official API provided by OpenAI (ope, 2023) and set the temperature to 0 to avoid randomness in token generation and ensure reproducibility. By default, we set the a top-$k$ token number to 3 and minimum cluster size to 100 in our hierarchical log sharder. For the LCU selector, we use an LCU sample size of 3, a interpolation factor $\lambda$ of 0.7, and a minimum similarity of 0.33, respectively. For parallelization, we distribute the jobs to 8 executors. We also conduct experiments of ULog with different LCU sample sizes and interpolation factors. As for baseline methods, we directly use the accuracy reported in previous work (Jiang et al., 2023b) and rerun them with their default parameters on the same environment to fairly compare the efficiency.
@@ 5. Evaluation Results
Table 1. Accuracy of ULog compared to state-of-the-art baselines on public datasets. (%)
@@ 5.1. RQ1: Effectiveness
Setup: In the first research question (RQ), we evaluate the accuracy of ULog by comparing it with state-of-the-art log parsers, as accuracy is the most critical factor for log parsers. For comprehensiveness, we compare ULog with both unsupervised parsers, which do not require human-labelled log templates, and label-required semantic-based log parsers, which rely on labelled log templates to support training, fine-tuning, or in-context learning. Table. 5 presents the overall results on four metrics for ULog compared to baseline methods. The best results for each metric on each dataset are highlighted in bold, while the second-best results are underlined. If a specific parser could not complete the parsing process within a reasonable timeframe (e.g., 12 hours), as per previous work (Zhu et al., 2019; Khan et al., 2022; Jiang et al., 2023a, b), we denote the score as “-”.
@@
Results: Overall, we observe that ULog achieves the best average grouping accuracy (GA), parsing accuracy (PA), and F1 score of template accuracy (FTA), as well as the second-best F1 score of grouping accuracy (FGA). Additionally, ULog demonstrates high accuracy across all 14 datasets, showcasing its robustness in parsing log data from diverse systems. For example, ULog achieves the highest FTA on 11 out of 14 datasets and the second-highest FTA on the remaining 3 datasets.
@@
Among unsupervised log parsers, it is evident that ULog significantly outperforms the other four baselines across all four evaluation metrics. Among the syntax-based parsers (i.e., AEL, Drain, and Brain), Brain achieves the highest scores in template-level metrics, with an average FGA of 71.8% and an average FTA of 35.4%. However, LUNAR exhibits a substantially higher average FGA of 91.4% and FTA of 81.6%, surpassing Brain by 19.6% and 46.2%, respectively. This is primarily due to the limitations of syntax-based parsers, which rely solely on manually crafted rules and thus struggle with complex and diverse log data. Furthermore, leveraging the extensive pre-trained knowledge of LLMs, LILAC w/o ICL demonstrates superior performance than syntax-based methods in PA, FGA, and FTA. However, our label-free LLM-based parser, ULog, significantly outperforms LILAC w/o ICL, with an average improvement of 15.2% in FGA and 26.9% in FTA, respectively. These substantial improvements underscore the effectiveness of ULog in harnessing the zero-shot capabilities of LLMs without labelled examples.
@@
Compared with supervised baseline parsers, ULog demonstrate comparable performance to existing state-of-the-art parser LILAC. Specifically, compared to LILAC, ULog has achieved superior performance in average GA (93.4%), PA (88.4%), and FTA (81.6%), while showing a slightly lower average FGA score (91.4%). LILAC retrieves similar logs and labelled templates to provide demonstrations during parsing, which are often inaccessible in real-world systems. However, ULog achieves a comparable performance to the state-of-the-art while does not need labelled templates, which is more robust and generalizable in practice. Additionally, compared to UniParser and LogPPT, ULog has achieved substantial improvements with an average improvements of 30.2% and 28.7% in all four metrics. This result again highlights the advantages of ULog, which is not also effective, but also requires no labelled data for training. The comparable performance of ULog indicates its potential to be applied in real-world production systems, which can address the label dependency problems of semantic-based log parsers as mentioned in Section 2.1.
@@ 5.2.1. Designs
In this research question, we first assess the individual contributions of the designs in the two components of ULog: the hierarchical log sharder and the two-stage LCU selector. To accomplish this, we have implemented several variants of ULog by either removing or replacing the designs in these components. Specifically, for the hierarchical log sharder, we created two variants by removing the respective clustering stages. Additionally, for the two-stage LCU selector, we replaced the hybrid LCU ranking algorithm with various alternatives: random selection, simple selection based on minimum or maximum similarity, and selection of consecutive log messages for querying the LLMs. To mitigate the impact of randomness, we also repeated experiment five times and calculated the mean scores as final results.
The average metric scores across all datasets are presented in Table 5.2.1. When the log length-based clustering is removed from the log sharder, the end-to-end accuracy of ULog drops from 93.4% to 87.7% and from 88.4% to 82.3%, respectively. Similar decreases are observed for the variant without top-k tokens clustering. These results indicate that both stages within the hierarchical log sharder significantly contribute to the overall performance of ULog. On the other hand, for the four variants related to the two-stage LCU selector, all four metrics show decreases across the board. For example, replacing the hybrid LCU ranking algorithm with random selection results in average GA and PA scores dropping by 7.4% and 6.6%, respectively. Notably, selecting the LCU based on maximum similarities among log messages has the most detrimental impact on ULog’s performance, resulting in a 10.2% decrease in GA and an 8.7% decrease in PA. This is primarily because selecting only the most similar log messages as an LCU for prompting prevents the LLM from accurately identifying templates and parameters by contrasting the tokens of these log messages. These results demonstrate that our proposed two-stage LCU selector is effective in selecting LCUs, thereby enabling LLMs to accurately parse log messages.
Table 2. Ablation study of components in ULog (%)
@@ 5.2.2. Configurations
In addition to the two key designs of ULog mentioned above, we have identified two configurations that could significantly affect ULog’s performance: the LCU sample size and the $\lambda$ values for LCU nomination. These configurations directly determine the LCUs for each query provided to the LLMs, and thus, they may significantly impact the accuracy of the LLMs’ parsed results. In this part, we explore different settings for the LCU sample size and the $\lambda$ values to evaluate their impact on ULog’s performance across all four evaluation metrics. Specifically, we first chose the default $\lambda$ value (i.e., 0.5) and varied the LCU sample sizes from 1 to 5. Additionally, we fixed the default LCU sample size (i.e., 3) and varied the $\lambda$ value from 0.0 to 1.0. The average metric scores under different settings across all datasets are presented in Figure 8.
Figure 8. Sensitivity study of configurations in ULog.
Based on the results shown in the left sub-figure of Figure 8, we observe that the performance of ULog remains consistently high across different LCU sample sizes. Notably, when the LCU sample size is set to 3, all four metrics achieve their highest scores. Conversely, smaller sample sizes result in lower accuracy. For instance, with a sample size of 1, the average GA and PA scores drop to approximately 89% and 86%, respectively, which is 4.5% and 2.5% lower than the scores achieved with a sample size of 3. This suggests that the absence of contrastive log messages can impair the parsing ability of LLMs. Interestingly, when the sample size is 2, the accuracy decreases further compared to a sample size of 1. This occurs because with only 2 contrastive log messages, the LLM tends to overfit to the differences between them, neglecting potential variable tokens in the log messages. On the other hand, increasing the LCU sample size to 4 or 5 results in a downward trend in accuracy. For example, the FTA score decreases from 81.6% to 80.6%. This indicates that larger LCU sample sizes may introduce additional noise and thus do not enhance the accuracy of ULog. Therefore, we have set the default LCU sample size to 3 in our experiments.
@@ 5.3. RQ3: Efficiency
Setup: Efficiency is an essential factor for log parsers in real-world usage, given the substantial volume of logs produced (Zhu et al., 2019; Wang et al., 2022). In this RQ, we evaluate the efficiency of ULog and all other baseline parsers by apply them in the large-scale datasets within Logpub, which contains average 3.6 million log messages per dataset. Specifically, we recorded the execution times for all log parsers in parsing all log datasets within Logpub, and then compute the average parsing time across all datasets. Since ULog is designed to enable parallelism for processing log buckets, we compute the average parsing time for both serial mode and parallel model of ULog. The efficiency results are demonstrated in the Figure 9.
Figure 9. Efficiency of ULog and baselines (second).
Results: The results demonstrate that ULog achieves higher efficiency compared to the most efficient semantic-based log parser, LILAC, and comparable efficiency to the most efficient syntax-based log parser, Drain. Specifically, the average parsing time for 3.6 million log messages in serial mode using ULog is 1191.4 seconds. In practice, we can leverage the parallel mode of ULog to achieve better efficiency, reducing the parsing time to just 532.9 seconds. For the most efficient syntax-based parser, Drain, which only takes an average of 488.5 seconds to parse each dataset, ULog achieve comparable efficiency, only slower by 9%. In contrast, the most efficient semantic-based parser, LILAC, takes an average of 620.1 seconds, which is 16.4% slower than ULog. Notably, even with GPU acceleration, other semantic-based parsers such as UniParser and LogPPT require significantly more time to parse large-scale log data. ULog outperforms them by 3.68 times and 6.75 times in efficiency, respectively. These results indicate that ULog is efficient in parsing large-scale log data, making it suitable for application in real-world production systems.

### v2

@@ 3.6. Parallelization
In previous sections, we have discussed the building components of LUNAR. In practice, scalability is a crucial issue when parsing vast amounts of log messages (Wang et al., 2022). Therefore, we introduce the parallelization mechanism of LUNAR to improve the parsing speed.
As mentioned in §3.2, hierarchical sharding produces mutually exclusive log buckets at two levels. Consequently, logs in different buckets can be processed in a completely parallel fashion. However, since bucket updating is performed among all the buckets of the same length at the first level. Simply allocating buckets at the second level to produce template and update buckets can lead to template conflict, i.e., different templates being assigned to the same log. To address the problem, the allocation is conducted at the first level. Specifically, we aim to allocate $n$ executors for all first-level log buckets. Within each first-level log bucket, LUNAR iteratively selects the largest second-level bucket, selects an LCU, queries the LLM to obtain the template, and updates buckets under this level.
@@ 4. Experiment Setup
We conduct extensive experiments to evaluate LUNAR by answering the following research questions (RQs):
  • RQ1: How effective is LUNAR?
  • RQ2: How do different settings affect LUNAR?
  • RQ3: How efficient is LUNAR?
@@ 4.1. Datasets
We evaluate LUNAR on Loghub-2.0 (He et al., 2020; Jiang et al., 2023b), a collection of large-scale benchmark datasets for log parsing provided by LogPAI team (Zhu et al., 2019). Loghub-2.0 is annotated with ground-truth templates for 14 diverse log datasets, which covers a variety of systems such as distributed systems, supercomputer systems, and server-side applications. Compared with the original Loghub (He et al., 2020), Loghub-2.0 is equipped with 3×3$\times$3 × templates and 1,800×1,800$\times$1 , 800 × log messages on average, which can support more comprehensive evaluation on accuracy and efficiency. On average, each dataset in Loghub-2.0 comprises 3.6 million log messages. In total, the collection contains approximately 3,500 unique log templates.
@@ 4.2. Baselines
We compare LUNAR with seven open-sourced state-of-the-art log parsers, including four unsupervised syntax-based parsers and three label-required semantic-based parsers. For unsupervised label-free parsers, we adopt three syntax-based methods and one LLM-based method. The syntax-based methods include AEL (Jiang et al., 2008), Drain (He et al., 2017), and Brain (Yu et al., 2023), chosen for their superior performance compared to other syntax-based methods (Zhu et al., 2019; Khan et al., 2022; Jiang et al., 2023b). As there are currently no LLM-based log parsers proposed for label-free parsing, we adopt a label-free variant of LILAC (Jiang et al., 2023a) (LILAC w/o ICL) by removing the in-context learning (ICL) module. In terms of label-required log parsers, we select three state-of-the-art semantic-based log parsers: UniParser (Liu et al., 2022), LogPPT (Le and Zhang, 2023), and LILAC (Jiang et al., 2023a). UniParser trains a long short-term memory (LSTM) model on labelled log data for log parsing. LogPPT uses labelled log data to perform prompt-based fine-tuning based on RoBERTa (Liu et al., 2019a). LILAC samples similar logs from a small set of labelled logs (e.g., 32 logs) for each log, utilizing an ICL paradigm for LLMs to parse logs.
@@ 4.3. Evaluation Metrics
Following previous works (Zhu et al., 2019; Khan et al., 2022; Jiang et al., 2023b), we evaluate LUNAR with the following four metrics.
@@ 4.4. Environment and Implementation
We conduct all the experiments on a Ubuntu 20.04.4 LTS server with 256RAM and an NVIDIA A100 40G GPU. The LLM used in LUNAR is ChatGPT (gpt-3.5-turbo-0613) due to its popularity in recent log analysis studies (Xu et al., 2024; Jiang et al., 2023a; Ma et al., 2024). We invoke the official API provided by OpenAI (ope, 2023) and set the temperature to 0 to avoid randomness in token generation and ensure reproducibility. By default, we set the a top-$k$ token number to 3 and minimum cluster size to 100 in our hierarchical log sharder. For the LCU selector, we use an LCU sample size of 3, a interpolation factor $\lambda$ of 0.7, and a minimum similarity of 0.33, respectively. For parallelization, we distribute the jobs to 8 executors. We also conduct experiments of LUNAR with different LCU sample sizes and interpolation factors. As for baseline methods, we directly use the accuracy reported in previous work (Jiang et al., 2023b) and rerun them with their default parameters on the same environment to fairly compare the efficiency.
@@ 5. Evaluation Results
Table 1. Accuracy of LUNAR compared to state-of-the-art baselines on public datasets. (%)
@@ 5.1. RQ1: Effectiveness
Setup: In the first research question (RQ), we evaluate the accuracy of LUNAR by comparing it with state-of-the-art log parsers, as accuracy is the most critical factor for log parsers. For comprehensiveness, we compare LUNAR with both unsupervised parsers, which do not require human-labelled log templates, and label-required semantic-based log parsers, which rely on labelled log templates to support training, fine-tuning, or in-context learning. Table. 5 presents the overall results on four metrics for LUNAR compared to baseline methods. The best results for each metric on each dataset are highlighted in bold, while the second-best results are underlined. If a specific parser could not complete the parsing process within a reasonable timeframe (e.g., 12 hours), as per previous work (Zhu et al., 2019; Khan et al., 2022; Jiang et al., 2023a, b), we denote the score as “-”.
@@
Results: Overall, we observe that LUNAR achieves the best average grouping accuracy (GA), parsing accuracy (PA), and F1 score of template accuracy (FTA), as well as the second-best F1 score of grouping accuracy (FGA). Additionally, LUNAR demonstrates high accuracy across all 14 datasets, showcasing its robustness in parsing log data from diverse systems. For example, LUNAR achieves the highest FTA on 11 out of 14 datasets and the second-highest FTA on the remaining 3 datasets.
@@
Among unsupervised log parsers, it is evident that LUNAR significantly outperforms the other four baselines across all four evaluation metrics. Among the syntax-based parsers (i.e., AEL, Drain, and Brain), Brain achieves the highest scores in template-level metrics, with an average FGA of 71.8% and an average FTA of 35.4%. However, LUNAR exhibits a substantially higher average FGA of 91.4% and FTA of 81.6%, surpassing Brain by 19.6% and 46.2%, respectively. This is primarily due to the limitations of syntax-based parsers, which rely solely on manually crafted rules and thus struggle with complex and diverse log data. Furthermore, leveraging the extensive pre-trained knowledge of LLMs, LILAC w/o ICL demonstrates superior performance than syntax-based methods in PA, FGA, and FTA. However, our label-free LLM-based parser, LUNAR, significantly outperforms LILAC w/o ICL, with an average improvement of 15.2% in FGA and 26.9% in FTA, respectively. These substantial improvements underscore the effectiveness of LUNAR in harnessing the zero-shot capabilities of LLMs without labelled examples.
@@
Compared with supervised baseline parsers, LUNAR demonstrate comparable performance to existing state-of-the-art parser LILAC. Specifically, compared to LILAC, LUNAR has achieved superior performance in average GA (93.3%), PA (88.4%), and FTA (81.6%), while showing a slightly lower average FGA score (91.4%). LILAC retrieves similar logs and labelled templates to provide demonstrations during parsing, which are often inaccessible in real-world systems. However, LUNAR achieves a comparable performance to the state-of-the-art while does not need labelled templates, which is more robust and generalizable in practice. Additionally, compared to UniParser and LogPPT, LUNAR has achieved substantial improvements with an average improvements of 30.2% and 28.7% in all four metrics. This result again highlights the advantages of LUNAR, which is not also effective, but also requires no labelled data for training. The comparable performance of LUNAR indicates its potential to be applied in real-world production systems, which can address the label dependency problems of semantic-based log parsers as mentioned in Section 2.1.
@@ 5.2.1. Designs
In this research question, we first assess the individual contributions of the designs in the two components of LUNAR: the hierarchical log sharder and the two-stage LCU selector. To accomplish this, we have implemented several variants of LUNAR by either removing or replacing the designs in these components. Specifically, for the hierarchical log sharder, we created two variants by removing the respective clustering stages. Additionally, for the two-stage LCU selector, we replaced the hybrid LCU ranking algorithm with various alternatives: random selection, simple selection based on minimum or maximum similarity, and selection of consecutive log messages for querying the LLMs. To mitigate the impact of randomness, we also repeated experiment five times and calculated the mean scores as final results.
The average metric scores across all datasets are presented in Table 5.2.1. When the log length-based clustering is removed from the log sharder, the end-to-end accuracy of LUNAR drops from 93.4% to 87.7% and from 88.4% to 82.3%, respectively. Similar decreases are observed for the variant without top-k tokens clustering. These results indicate that both stages within the hierarchical log sharder significantly contribute to the overall performance of LUNAR. On the other hand, for the four variants related to the two-stage LCU selector, all four metrics show decreases across the board. For example, replacing the hybrid LCU ranking algorithm with random selection results in average GA and PA scores dropping by 7.4% and 6.6%, respectively. Notably, selecting the LCU based on maximum similarities among log messages has the most detrimental impact on LUNAR’s performance, resulting in a 10.2% decrease in GA and an 8.7% decrease in PA. This is primarily because selecting only the most similar log messages as an LCU for prompting prevents the LLM from accurately identifying templates and parameters by contrasting the tokens of these log messages. These results demonstrate that our proposed two-stage LCU selector is effective in selecting LCUs, thereby enabling LLMs to accurately parse log messages.
Table 2. Ablation study of components in LUNAR (%)
@@ 5.2.2. Configurations
In addition to the two key designs of LUNAR mentioned above, we have identified two configurations that could significantly affect LUNAR’s performance: the LCU sample size and the $\lambda$ values for LCU nomination. These configurations directly determine the LCUs for each query provided to the LLMs, and thus, they may significantly impact the accuracy of the LLMs’ parsed results. In this part, we explore different settings for the LCU sample size and the $\lambda$ values to evaluate their impact on LUNAR’s performance across all four evaluation metrics. Specifically, we first chose the default $\lambda$ value (i.e., 0.5) and varied the LCU sample sizes from 1 to 5. Additionally, we fixed the default LCU sample size (i.e., 3) and varied the $\lambda$ value from 0.0 to 1.0. The average metric scores under different settings across all datasets are presented in Figure 8.
Figure 8. Sensitivity study of configurations in LUNAR.
Based on the results shown in the left sub-figure of Figure 8, we observe that the performance of LUNAR remains consistently high across different LCU sample sizes. Notably, when the LCU sample size is set to 3, all four metrics achieve their highest scores. Conversely, smaller sample sizes result in lower accuracy. For instance, with a sample size of 1, the average GA and PA scores drop to approximately 89% and 86%, respectively, which is 4.5% and 2.5% lower than the scores achieved with a sample size of 3. This suggests that the absence of contrastive log messages can impair the parsing ability of LLMs. Interestingly, when the sample size is 2, the accuracy decreases further compared to a sample size of 1. This occurs because with only 2 contrastive log messages, the LLM tends to overfit to the differences between them, neglecting potential variable tokens in the log messages. On the other hand, increasing the LCU sample size to 4 or 5 results in a downward trend in accuracy. For example, the FTA score decreases from 81.6% to 80.6%. This indicates that larger LCU sample sizes may introduce additional noise and thus do not enhance the accuracy of LUNAR. Therefore, we have set the default LCU sample size to 3 in our experiments.
@@ 5.3. RQ3: Efficiency
Setup: Efficiency is an essential factor for log parsers in real-world usage, given the substantial volume of logs produced (Zhu et al., 2019; Wang et al., 2022). In this RQ, we evaluate the efficiency of LUNAR and all other baseline parsers by apply them in the large-scale datasets within Logpub, which contains average 3.6 million log messages per dataset. Specifically, we recorded the execution times for all log parsers in parsing all log datasets within Logpub, and then compute the average parsing time across all datasets. Since LUNAR is designed to enable parallelism for processing log buckets, we compute the average parsing time for both serial mode and parallel mode of LUNAR. The efficiency results are demonstrated in the Figure 9.
Figure 9. Efficiency of LUNAR and baselines (second).
Results: The results demonstrate that LUNAR achieves higher efficiency compared to the most efficient semantic-based log parser, LILAC, and comparable efficiency to the most efficient syntax-based log parser, Drain. Specifically, the average parsing time for 3.6 million log messages in serial mode using LUNAR is 1191.4 seconds. In practice, we can leverage the parallel mode of LUNAR to achieve better efficiency, reducing the parsing time to just 532.9 seconds. For the most efficient syntax-based parser, Drain, which only takes an average of 488.5 seconds to parse each dataset, LUNAR achieve comparable efficiency, only slower by 9%. In contrast, the most efficient semantic-based parser, LILAC, takes an average of 620.1 seconds, which is 16.4% slower than LUNAR. Notably, even with GPU acceleration, other semantic-based parsers such as UniParser and LogPPT require significantly more time to parse large-scale log data. LUNAR outperforms them by 3.68 times and 6.75 times in efficiency, respectively. These results indicate that LUNAR is efficient in parsing large-scale log data, making it suitable for application in real-world production systems.